# LMSeg: A deep graph message-passing network for efficient and accurate semantic segmentation of large-scale 3D landscape meshes

### v1

@@ 1 Introduction
Recent advancements in data acquisition and terrain reconstruction methods enabled efficient, automated modelling of large-scale landscape terrains. Modelling of landscapes relies on the representation of the terrain and potential spatial objects on the terrain as 2.5D rasters (regularly-spaced cells with elevation), or as textured 3D meshes of irregular triangles. Compared to 3D point cloud or 2.5D raster landscape representations, 3D mesh modelling of landscapes explicitly capture geometric information and semantic attributes (e.g., texture/landcover, aspect, gradient) associated with triangle surfaces encoding explicit face adjacency. Current 3D mesh segmentation methods are, however, still underdeveloped and underused compared to raster segmentation, and are facing significant challenges with: (1) the accuracy of segmentation, especially on small objects; (2) the tractability and efficiency of computation (i.e, efficient model training and inference); and with (3) evaluation on diverse, large-scale landscape mesh datasets, as discussed in a recent survey (Adam et al., 2023). These challenges explicitly point out the ongoing problems in existing landscape segmentation models, suggesting limited accuracy in segmenting complex terrain structures and high computational complexity that limits the model applications to large-scale landscape mesh datasets (Kölle et al., 2021; Gao et al., 2021). With the segmentation accuracy and efficiency issues, we pose the following question: How to improve the accuracy and efficiency of mesh segmentation by leveraging the geometric (i.e., face normals and Euclidean coordinates) and semantic (i.e., RGB or HSV color) information of large-scale landscape meshes? We hypothesize that the barycentric dual graph representation of landscape meshes can tightly capture the rich semantic and informative topography of mesh surfaces, which enable end-to-end model trainability with graph-based learning architectures (Gilmer et al., 2017; Battaglia et al., 2018) and facilitate expressive feature learning for accurate and efficient landscape segmentation.
@@ 2.3 Learning Models for Urban Meshes
To evaluate mesh segmentation and scene understanding methods on real-world landscape mesh data, a learning-based approach by Tutzauer et al. (2019) computes the barycenter of landscape surface mesh triangles as point clouds. A multi-branch 1D CNN network is trained on the geometric and radiometric features of point clouds (i.e., the barycenter of landscape meshes) to predict per-face labels. Later, Laupheimer et al. (2020) associated 3D lidar point clouds with textured meshes and apply a PointNet++ (Qi et al., 2017b) for mesh segmentation. The method by Tang et al. (2022) abstracts the triangular meshes into the barycentric point clouds of meshes and applies a point-based Transformer architecture (Zhao et al., 2021) for downstream mesh segmentation. However, the point-based transformer learning architecture (Zhao et al., 2021) in Tang et al. (2022) is comparatively computationally expensive and requires over 40 hours of training time on a proprietary urban dataset for mesh segmentation, which significantly limits the model applicability for larger mesh datasets. The approach proposed by Yang et al. (2023) represents the triangle meshes of landscape as surface graphs (i.e., barycentric dual graphs) and then applies a hierarchical deep network for mesh segmentation. This approach resembles a U-Net segmentation architecture (Ronneberger et al., 2015) and hierarchically constructs surface connective neighbourhoods (SC neighborhood) of barycentric dual graphs for hierarchical feature learning on input meshes. Additional TextureConv blocks based on 2D convolution networks are required to learn facet texture features on the textures of triangle meshes for semantic segmentation but limits the end-to-end trainability of proposed model. PSSNet by Weixiao et al. (2023) proposed a two-stage deep learning-based architecture for urban mesh segmentation, where the geometric node and edge features are extracted by deep neural networks for superfacet classification. The two-stage PSSNet learning architecture requires a super-point grouping and edge dependency extraction process in the first stage for the subsequent model learning in the second stage. This model design fails to jointly optimise the two-stage model architecture and therefore is prohibitive to end-to-end model trainability and causing inefficient model inference.
@@ 4.4 Results
We present LMSeg validation results (on Area 4) and test results (on Area 2) in Table 2 and the qualitative performance in Fig. 7. For comparison, we present the test performance of point-based learning methods (i.e., RandLANet (Hu et al., 2020), PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021)) and graph-based learning methods (i.e., DeeperGCNs (Li et al., 2023) and Graph U-Net (Gao and Ji, 2019)). Note that we adopt different point-based learning methods and additional graph-based methods for benchmarking of the BBW dataset to restrict the model size and training parameters to a comparable size to LMSeg.
For fair evaluation, the graph-based methods tested on BBW dataset take the barycentric graph of the processed landscape mesh as inputs whereas the point-based methods take the vertex points of the barycentric graph as inputs. Overall, LMSeg achieves the best mIoU of 59.5% and 58.5%, F1 score of 74.6% and 74.0% among all test methods, on test (Area 2) and validation (Area 4) sets.
Different to urban landscape meshes in SUM (Fig. 6), the natural landscape meshes (Fig. 7) in BBW have dense, near-uniformly distributed triangles capturing the fine-grained geometric objects (i.e., walls and ditches) on the terrain surface. Point-based learning methods (i.e., RandLANet (Hu et al., 2020): 50.4% and 53.7% mIoU, Point Transformer (Zhao et al., 2021): 42.6% and 47.5% mIoU, PointNet++ (Qi et al., 2017b): 42.9% and 43.3% mIoU) perform generally better than graph-based learning methods with deep residual graph convolution networks (i.e., DeeperGCNs (Li et al., 2023): 32.8% and 35.5% mIoU and Graph U-Net (Gao and Ji, 2019): 25.1% and 28.9 mIoU) operating on the barycentric graph of meshes.
@@
The error maps of LMSeg on BBW dataset (Fig. 4) show that the prediction errors mostly appear at the ambiguous object boundaries between the irregular stone wall structures and the remainder of the terrain surface. This is where mesh faces might be impacted by misclassified labels from human annotators.
@@
SUM-A Ablated model without HGA+ components.
SUM-B Ablated model without LGA+ components.
SUM-C Ablated model without learnable aggregation.
SUM-D and E Ablated models with $8$- and $12$-nbrs.
SUM-F Ablated model without SUM HSV colour.
SUM-G Ablated model without SUM face normals.
BBW-A and B Ablated models tested on 75%-res. meshes.
BBW-C and D Ablated models tested on 50%-res. meshes.
BBW-E and F Ablated models tested on 30%-res. meshes.
BBW-G and H Ablated models without BBW HSV colour.
BBW-I and J Ablated models without BBW face normals.
@@ 5 Discussion
SUM-D and E study the dependence of the model performance on the size of $k$-nearest neighbours in HGA+ layers of LMSeg. Empirical results show that reducing $k$-nearest neighbours to smaller values leads to fluctuating mIoU results: 73.0% (LMSeg, $k$-nbrs: 16) $\rightarrow$ 73.1% (SUM-D, $k$-nbrs: 12) and 68.1% (SUM-E, $k$-nbrs: 8). The nearest neighbours $k$ is therefore a potential hyper-parameter that can be fine-tuned for the target mesh dataset. Potentially, increasing $k$ to larger values enables capturing richer geometric information on hierarchical sub-sampled barycentric graph $\mathcal{G}^{l-1}(\mathcal{M})\rightarrow\mathcal{G}^{l}(\mathcal{M})$ for better model performances but comes at high computing cost.
@@
We investigate LMSeg’s performance in function of mesh resolution (i.e., natural landscape meshes at 75%, 50% and 30% resolutions) on BBW dataset (Table 5). The meshes are simplified for evaluation by QEM simplification (Garland and Heckbert, 1997) to coarser mesh resolutions and converted to barycentric graphs as model inputs. At 75% mesh resolutions, models BBW-A (58.5% $\rightarrow$ 59.1% on Area 4) and BBW-E (59.5% $\rightarrow$ 60.6% on Area 2) show slight model improvements of mIoU compared to LMSeg. At 50% mesh resolutions, models BBW-B (58.5% $\rightarrow$ 53.4% on Area 4) and BBW-F (59.5% $\rightarrow$ 56.3% on Area 2) suffer minor degradation of mIoU. At 30% mesh resolutions, models BBW-C (58.5% $\rightarrow$ 37.4% on Area 4) and BBW-G (59.5% $\rightarrow$ 42.2% on Area 2) exhibit significant degradation of mIoU. These empirical results demonstrate LMSeg’s robustness on natural landscapes of changing resolutions.
@@
Models BBW-G, H and BBW-I, J (Table 3) test the impact of HSV colour and face normals, respectively, on the BBW validation and test sets. For models BBW-G and H, the ablation of HSV colour features exhibits relatively small test and validation performance drops (58.5% $\rightarrow$ 54.7% on Area 4 and 59.5% $\rightarrow$ 56.4% on Area 2). For models BBW-I and J, the ablation of face normals shows negligible performance change (58.5% $\rightarrow$ 58.3% on Area 4 and 59.5% $\rightarrow$ 59.4% on Area 2). This could be attributed to the fact that, the geometric mesh features (i.e., face normals) of Terra. and Wall classes are both pointing upward but at varying angles, and the semantic information of stone walls is largely perturbed by bush/tree objects as shown in Fig. 7, rendering these mesh features less effective for object segmentation.
@@ 6 Conclusions
In this study, we propose LMSeg, an efficient deep graph message-passing network for accurate semantic segmentation of large-scale 3D landscape meshes. The LMSeg model takes the barycentric dual graphs of landscape meshes as inputs, enabling end-to-end model trainability with deep graph message-passing networks. It fully leverages the geometric and texture features of triangular surface meshes (i.e., face normals and color information) for accurate landscape mesh segmentation. LMSeg effectively learns hierarchical and local geometric features from the barycentric dual graphs of meshes, thanks to the learnable feature aggregation in Geometry Aggregation Plus (GA+) layers.
Our empirical results on large-scale urban and natural landscape datasets suggest that LMSeg outperforms existing mesh segmentation methods in both accuracy and model inference efficiency. The model shows notable improvement in segmenting small mesh objects (e.g., vehicles and boats) in urban landscapes and irregular stone wall structures in natural landscapes.
However, the proposed LMSeg exhibits somewhat lower segmentation performance on mesh objects with geometrically similar barycentric graph structures (e.g., water and terrain in SUM) due to limited network receptive fields. Future research will focus on developing a deep graph message-passing network capable of learning more fine-grained semantic mesh features and capturing comprehensive and global spatial information of landscape meshes with a larger receptive field.
@@ Acknowledgements
We thank Gunditj Mirring Traditional Owners Corporation (GMTOC) for providing the research data and overseeing the usage of research data in this study. Their cooperation and support have been invaluable. We also thank Mashnoon Islam for his dedicated data annotation work for this study. His meticulous efforts significantly contributed to the quality and accuracy of our research.

### v2

@@ 1 Introduction
Recent advancements in data acquisition and terrain reconstruction methods enabled efficient, automated modelling of large-scale landscape terrains. Modelling of landscapes relies on the representation of the terrain and potential spatial objects on the terrain as 2.5D rasters (regularly-spaced cells with elevation), or as textured 3D meshes of irregular triangles. Compared to 3D point cloud or 2.5D raster landscape representations, 3D mesh modelling of landscapes explicitly capture geometric information and semantic attributes (e.g., texture/landcover, aspect, gradient) associated with triangle surfaces encoding explicit face adjacency. Current 3D mesh segmentation methods are, however, still underdeveloped and underused compared to raster segmentation, and are facing significant challenges with: (1) the accuracy of segmentation, especially on small objects; (2) the tractability and efficiency of computation (i.e, efficient model training and inference); and with (3) evaluation on diverse, large-scale landscape mesh datasets, as discussed in a recent survey (Adam et al., 2023). These challenges explicitly point out the ongoing problems in existing landscape segmentation models, suggesting limited accuracy in segmenting complex terrain structures and high computational complexity that limits the model applications to large-scale landscape mesh datasets (Kölle et al., 2021; Gao et al., 2021). With the segmentation accuracy and efficiency issues, we pose the following question: How to improve the accuracy and efficiency of mesh segmentation by leveraging the geometric (i.e., face normals and Euclidean coordinates) and semantic (i.e., RGB or HSV color) information of large-scale landscape meshes? We hypothesize that the barycentric dual graph representation of landscape meshes can tightly capture the rich semantic and informative topography of mesh surfaces, which enable end-to-end model trainability with graph-based learning architectures (Gilmer et al., 2017) and facilitate expressive feature learning for accurate and efficient landscape segmentation.
@@ 2.3 Learning Models for Urban Meshes
To evaluate mesh segmentation and scene understanding methods on real-world landscape mesh data, a learning-based approach by Tutzauer et al. (2019) computes the barycenter of landscape surface mesh triangles as point clouds. A multi-branch 1D CNN network is trained on the geometric and radiometric features of point clouds (i.e., the barycenter of landscape meshes) to predict per-face labels. Later, Laupheimer et al. (2020) associated 3D lidar point clouds with textured meshes and apply a PointNet++ (Qi et al., 2017b) for mesh segmentation. The method by Tang et al. (2022) abstracts the triangular meshes into the barycentric point clouds of meshes and applies a point-based Transformer architecture (Zhao et al., 2021) for downstream mesh segmentation. However, the point-based transformer learning architecture (Zhao et al., 2021) in Tang et al. (2022) is comparatively computationally expensive and requires over 40 hours of training time on a proprietary urban dataset for mesh segmentation, which significantly limits the model applicability for larger mesh datasets. The approach proposed by Yang et al. (2023) represents the triangle meshes of landscape as surface graphs (i.e., barycentric dual graphs) and then applies a hierarchical deep network for mesh segmentation. This approach resembles a point-based segmentation architecture (Qi et al., 2017b) and hierarchically constructs surface connective neighbourhoods (SC neighborhood) of barycentric dual graphs for hierarchical feature learning on input meshes. Additional TextureConv blocks based on 2D convolution networks are required to learn facet texture features on the textures of triangle meshes for semantic segmentation but limits the end-to-end trainability of proposed model. PSSNet by Weixiao et al. (2023) proposed a two-stage deep learning-based architecture for urban mesh segmentation, where the geometric node and edge features are extracted by deep neural networks for superfacet classification. The two-stage PSSNet learning architecture requires a super-point grouping and edge dependency extraction process in the first stage for the subsequent model learning in the second stage. This model design fails to jointly optimise the two-stage model architecture and therefore is prohibitive to end-to-end model trainability and causing inefficient model inference.
@@ 4.4 Results
We present LMSeg validation results (on Area 4) and test results (on Area 2) in Table 2 and the qualitative performance in Fig. 7. For comparison, we include different point-based learning methods (i.e., RandLANet (Hu et al., 2020), PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021)) and graph-based learning methods (i.e., DeeperGCNs (Li et al., 2023) and Graph U-Net (Gao and Ji, 2019)), ensuring comparable model sizes and training parameters. In practice, the graph-based methods take the barycentric graph of the processed landscape mesh as inputs and the point-based methods take the vertex points of the barycentric graph as inputs. Overall, LMSeg achieves the best mIoU of 59.5% and 58.5%, F1 score of 74.6% and 74.0% among all methods on test (Area 2) and validation (Area 4) sets.
Natural landscape meshes in BBW, unlike urban landscape meshes in SUM, have dense, uniformly distributed triangles capturing fine-grained geometric objects (i.e., walls and ditches) of the terrain surface. Point-based learning methods (i.e., RandLANet (Hu et al., 2020): 50.4% and 53.7% mIoU, Point Transformer (Zhao et al., 2021): 42.6% and 47.5% mIoU, PointNet++ (Qi et al., 2017b): 42.9% and 43.3% mIoU) perform generally better than graph-based learning methods with deep residual graph convolution networks (i.e., DeeperGCNs (Li et al., 2023): 32.8% and 35.5% mIoU and Graph U-Net (Gao and Ji, 2019): 25.1% and 28.9 mIoU) operating on the barycentric graph of meshes.
@@
Error maps of LMSeg on the BBW dataset (Fig. 4) show prediction errors mainly at ambiguous boundaries between irregular stone walls and the terrain, likely due to misclassified labels from human annotators.
@@
SUM-A Without HGA+ components.
SUM-B Without LGA+ components.
SUM-C Without learnable feature aggregation.
SUM-D and E With $8$- and $12$-nearest neighbours.
SUM-F Without SUM HSV colour.
SUM-G Without SUM face normals.
BBW-A and B Tested on 75%-res. meshes.
BBW-C and D Tested on 50%-res. meshes.
BBW-E and F Tested on 30%-res. meshes.
BBW-G and H Without BBW HSV colour.
BBW-I and J Without BBW face normals.
@@ 5 Discussion
SUM-D and E examine the effect of $k$-nearest neighbors in HGA+ layers on model performance. Results show that reducing $k$ leads to fluctuating mIoU: 73.0% (LMSeg, $k=16$) $\rightarrow$ 73.1% (SUM-D, $k=12$) and 68.1% (SUM-E, $k=8$). Therefore, $k$ is a tunable hyper-parameter for the target mesh dataset. Increasing $k$ can potentially capture richer geometric information on hierarchical sub-sampled barycentric graph $\mathcal{G}^{l-1}(\mathcal{M})\rightarrow\mathcal{G}^{l}(\mathcal{M})$ for better performance but at a higher computational cost.
@@
We investigated LMSeg’s performance on natural landscape meshes at 75%, 50%, and 30% resolutions of the BBW dataset (Table 5). The meshes were simplified via QEM simplification to coarser resolutions and converted to barycentric graphs for input. At 75% resolution, models BBW-A (58.5% $\rightarrow$ 59.1% on Area 4) and BBW-E (59.5% $\rightarrow$ 60.6% on Area 2) showed slight mIoU improvements. At 50% resolution, models BBW-B (58.5% $\rightarrow$ 53.4% on Area 4) and BBW-F (59.5% $\rightarrow$ 56.3% on Area 2) suffered minor mIoU degradation. At 30% resolution, models BBW-C (58.5% $\rightarrow$ 37.4% on Area 4) and BBW-G (59.5% $\rightarrow$ 42.2% on Area 2) experienced significant mIoU degradation. These results demonstrate LMSeg’s robustness across varying landscape resolutions.
@@
Models BBW-G, H, and BBW-I, J (Table 3) assessed the impact of HSV color and face normals on the BBW validation and test sets. Removing HSV color features (i.e., BBW-G and H) resulted in relatively small performance drops (58.5% $\rightarrow$ 54.7% on Area 4 and 59.5% $\rightarrow$ 56.4% on Area 2). The removal of face normals ((i.e., BBW-I and J) showed negligible changes (58.5% $\rightarrow$ 58.3% on Area 4 and 59.5% $\rightarrow$ 59.4% on Area 2). This is likely because the geometric features (face normals) of terrain and wall classes are similar, and the semantic information of stone walls is affected by bush/tree objects, as shown in Fig. 7, making these features less effective for segmentation.
@@ 6 Conclusions
In this study, we introduce LMSeg, an efficient deep graph message-passing network for accurate semantic segmentation of large-scale 3D landscape meshes. LMSeg utilizes barycentric dual graphs of landscape meshes, enabling end-to-end model training with deep graph message-passing networks. It leverages geometric and texture features of triangular surface meshes (face normals and color information) for accurate segmentation, learning hierarchical and local geometric features through Geometry Aggregation Plus (GA+) layers.
Our results on large-scale urban and natural landscape datasets show that LMSeg outperforms existing mesh segmentation methods in accuracy and inference efficiency, with notable improvements in segmenting small mesh objects (i.e., Vehic. and Boat in SUM) and irregular stone wall structures.
However, LMSeg shows lower performance on mesh objects with geometrically similar barycentric graph structures (e.g., Water and Terra. in SUM) due to limited network receptive fields. Future research will focus on developing a network capable of learning finer semantic mesh features and capturing comprehensive global spatial information with a larger receptive field.
@@ Acknowledgements
We thank Gunditj Mirring Traditional Owners Corporation (GMTOC) for providing the research data and overseeing its usage in this study, as well as Mashnoon Islam for his dedicated data annotation work; their invaluable cooperation, support, and meticulous efforts have significantly contributed to the quality of our research.